# Reservoir Predictive Path Integral Control for Unknown Nonlinear Dynamics

### v1

@@ IV-A Echo State Network
We describe the ESN model and its learning method. We employ the ESN with leaky terms, which explicitly incorporates temporal dynamics through leak rates of the target control system [20]. For the learning method, we employ online learning via RLS [35].
IV-A1 Model
The input-output relationship in 1 and 2 is modeled using the following ESN [20]:
${\mathbf{x}}_{t+1}$ $=(1-\alpha){\mathbf{x}}_{t}+\alpha\,f\left({\mathbf{W}}^{\text{res}}{\mathbf{x}}_{t}+{\mathbf{W}}^{\text{in }}{\mathbf{u}}_{t}\right),$ (6)
@@ IV-A1 Model
Here, $^{\mathbf{x}}_{t}\in{\mathbb{R}}^{\hat{N}}$ is the reservoir state vector, where $\hat{N}\in{\mathbb{N}}$, and $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{L}$ is the output vector. Additionally, $^{\mathbf{W}}^{\text{res}}\in{\mathbb{R}}^{\hat{N}\times\hat{N}}$ and ${\mathbf{W}}^{\text{in}}\in{\mathbb{R}}^{\hat{N}\times M}$ are the state and input weight matrices, respectively, and ${\mathbf{W}}^{\text{out}}_{t}\in{\mathbb{R}}^{\hat{N}\times L}$ is the output weight matrix. Furthermore, $f\colon{\mathbb{R}}^{\hat{N}}\to{\mathbb{R}}^{\hat{N}}$ is an element-wise nonlinear activation function called the activation function, and $\alpha\in(0,1]$ is a parameter called the leak rate.
In contrast to conventional NNs, the state and input weight matrices ${\mathbf{W}}^{\text{res}}$ and ${\mathbf{W}}^{\text{in}}$ remain fixed as randomly initialized matrices. In many cases, the reservoir state dimension $\hat{N}$ is set greater than the input or output dimensions. The matrix ${\mathbf{W}}^{\text{res}}$ is a sparse random matrix with many components equal to zero, designed empirically so that the maximum absolute value of eigenvalues (spectral radius) is 1 or less. The matrix ${\mathbf{W}}^{\text{in}}$ is randomly generated from a uniform distribution. The reservoir state dimension $\hat{N}$, sparseness and spectral radius of ${\mathbf{W}}^{\text{res}}$, the shape of the activation function $f$, the leak rate $\alpha$, and the range of the input weight matrix ${\mathbf{W}}^{\text{in}}$ are hyperparameters that must be set prior to learning.
IV-A2 Learning Algorithm
Given : Discount factor $\gamma$, previous output weight ${\mathbf{W}}^{\mathrm{out}}_{t-1}$, precision matrix ${\mathbf{P}}_{t-1}$, observed output ${\mathbf{y}}_{t}$, and current state ${\mathbf{x}}_{t}$.
return ${\mathbf{W}}^{\mathrm{out}}_{t}$ and ${\mathbf{P}}_{t}$
Algorithm 2 RLS [35]
We describe the learning method for the output weight matrix ${\mathbf{W}}^{\text{out}}$, which is the only learnable parameter in ESN [35]. This weight matrix is learned to minimize the error between the observed output and the model output, written as follows:
@@ IV-A2 Learning Algorithm
where $\gamma\in(0,1]$ is the discount factor.
@@
The sequential learning method that repeatedly applies 9 and 10 with arbitrarily initialized parameters ${\mathbf{W}}_{-1}^{\text{out}}$ and ${\mathbf{P}}_{-1}$ is called RLS [35]. The RLS algorithm for ESN is shown in Algorithm 2.
The following proposition states that the precision matrix ${\mathbf{P}}_{t}$ of the variable ${\mathbf{x}}_{t}$ represents the covariance of the estimated output weight matrix ${\mathbf{W}}^{\text{out}}_{t}$. This result is obtained by extending the analysis in Ref. [35] to the case of ESNs, where the estimated parameters take a matrix form.
@@
Assume that there exists a weight $\bar{\mathbf{W}}^{\mathrm{out}}\in{\mathbb{R}}^{\hat{N}\times L}$ such that at any time $t$, the relationship between the reservoir state and the true plant output is expressed as
${\mathbf{y}}_{t}=\bar{\mathbf{W}}^{\text{out}\top}{\mathbf{x}}_{t}+\mbox{\boldmath$\epsilon$}_{t},$ (11)
where $\mbox{\boldmath$\epsilon$}_{t}\in{\mathbb{R}}^{L}$ is noise that follows a normal distribution with mean 0 and covariance matrix $\tilde{\Sigma}=\operatorname*{diag}(\sigma_{1}^{2},\ldots,\sigma_{L}^{2})$ independently at each time. Then, if $\gamma=1$,
${\mathbb{E}}[\mathrm{vec}({\mathbf{W}}^{\mathrm{out}*}_{t}-\bar{\mathbf{W}}^{\mathrm{out}})\mathrm{vec}({\mathbf{W}}^{\mathrm{out}*}_{t}-\bar{\mathbf{W}}^{\mathrm{out}})^{\top}]=\tilde{\Sigma}\otimes{\mathbf{P}}_{t}.$ (12)
This property is used when adding perturbations to the output weight matrix in UMPPI, as described in Section IV-B.
@@ IV-B Uncertainty-aware Model Predictive Path Integral Control
We propose control methods combining the ESN with MPPI or its uncertainty-aware variant UMPPI. In MPPI, multiple noisy input sequences are sampled based on a reference input to simulate future trajectories of the control system, and an output cost is computed for each sequence [27]. The optimal control input is approximately computed as a weighted average of many samples, where samples with lower costs receive higher weights. In UMPPI, in addition to sampling input sequences, output weight matrices of the ESN with noise added are sampled. The magnitude of the noise applied to the output matrices is determined by the precision matrix ${\mathbf{P}}_{t}$ in RLS. This provides control input perturbations according to the identification accuracy, resulting in automatic adjustment of exploration and exploitation. Note that in Algorithm 3 described later below, when the sample size ${\tilde{K}}=1$ and the noise covariance $\tilde{\Sigma}=0$, UMPPI reduces to the standard MPPI algorithm. Therefore, we focus on the UMPPI formulation in the following description.
@@
The following derivation of UMPPI extends the framework of Ref. [27] by introducing perturbations of the output-weight matrix. This extension preserves the mathematical consistency of the MPPI formulation and constitutes a main contribution of the present work. Specifically, Proposition IV.4 parallels Eq. (13) in Ref. [27]; Proposition IV.5 is established by the same argument as in Section III-A of that reference; and Proposition IV.6 corresponds to Eq. (27) in the same reference.
To explain the theoretical foundation of the proposed control algorithm, we outline how UMPPI computes the optimal control solution through a sequence of sampling-based approximations. The derivation of UMPPI proceeds in the following three steps:
@@
    Stochastic reformulation. The original MPC problem is first recast as a stochastic optimization over noise-injected input sequences, as expressed in 20.
@@
    Free-energy relaxation. Introducing a free-energy functional converts this stochastic problem into the minimization of a Kullback–Leibler divergence between two probability distributions (see 31).
@@
This formulation offers a sampling-based interpretation of MPC, where control sequences are drawn from a distribution and weighted according to their associated costs. The details of each step are described in Sections IV-B1, IV-B2 and IV-B3.
@@ IV-B1 Stochastic reformulation
This subsection reformulates the original MPC problem in 3 as a stochastic optimization problem over perturbed control inputs and model parameters in 20. Like conventional MPPI, UMPPI considers perturbed inputs ${\mathbf{v}}_{t}\sim{\mathcal{N}}({\mathbf{u}}_{t},\Sigma)$ where noise is applied to some initial inputs ${\mathbf{u}}_{t}$ as follows:
$\hat{\mathbf{x}}_{\tau+1}=\hat{\mathbf{F}}(\hat{\mathbf{x}}_{\tau},{\mathbf{v}}_{\tau}),\ \hat{\mathbf{x}}_{t}={\mathbf{x}}_{t}.$ (13)
Here, $\hat{\mathbf{F}}\colon{\mathbb{R}}^{\hat{N}}\times{\mathbb{R}}^{M}\to{\mathbb{R}}^{\hat{N}}$ represents the state update equation in 6:
$\hat{\mathbf{F}}(\hat{\mathbf{x}}_{\tau},{\mathbf{v}}_{\tau})\coloneq(1-\alpha)\hat{\mathbf{x}}_{\tau}+\alpha f\left({\mathbf{W}}^{\text{res}}\hat{\mathbf{x}}_{\tau}+{\mathbf{W}}^{\text{in }}{\mathbf{v}}_{\tau}\right).$ (14)
Furthermore, $\Sigma\in{\mathbb{R}}^{M\times M}$ represents the covariance matrix of the applied noise, and its value is determined based on the input weight matrix $R$ of the cost function (3) as follows:
@@
where $\lambda\in{\mathbb{R}}$ is a parameter called the inverse temperature, which is a design parameter. In addition to this, UMPPI considers outputs using weight matrices ${\mathbf{W}}_{\tau}$ with noise applied to the reference output weight matrix ${\mathbf{W}}^{\mathrm{out}}_{\tau}$ at each time, which are expressed as
@@
where $\mathrm{vec}({\mathbf{W}}_{\tau})\sim{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}^{\mathrm{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t})$, which is justified by Proposition IV.2. Applying noise to inputs enables approximate computation of optimal inputs through the relationship with free energy. Additionally, applying noise to the output weight matrix provides control input perturbations according to identification accuracy at each time, resulting in automatic adjustment of exploration and exploitation by the controller.
@@
$\begin{aligned} &p(W_{t}|{\mathbf{W}}_{t}^{\text{out}},{\mathbf{P}}_{t},\tilde{\Sigma})\\ &\quad=\prod_{\tau=t}^{t+H}{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}_{\tau})|\mathrm{vec}({\mathbf{W}}^{\mathrm{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}).\end{aligned}$ (19)
@@ IV-B2 Free energy relaxation
This subsection introduces a variational relaxation of the stochastic optimization in 20 using the concept of free energy, leading to an alternative minimization problem in 31. We define the free energy $\mathcal{F}$} as follows:
@@
$q(W_{t})$ $\coloneqq\prod_{\tau=t}^{t+H}{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}_{\tau})|\mathrm{vec}({\mathbf{W}}^{\mathrm{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}).$ (26)
Here, ${\mathbf{u}}^{\text{ref}}$ is the reference input defined in 4. The following two propositions show the relationship between LABEL:eq:free_energy and 20.
@@
Suppose that $V_{t}$ and $W_{t}$ in 18 and LABEL:eq:prob_W follow the distributions given by
@@
Then, equality holds for 27:
@@ IV-B3 Sampling-based optimization
Suppose that the admissible control input set is given by ${\mathcal{U}}={\mathbb{R}}^{M\times H}$. Then, given any input sequence $\hat{U}_{t}$, the control input satisfying 31 is computed as
@@
Given : Sample size of input/output perturbation $K,{\tilde{K}}$, prediction horizon $H$, initial estimate of input sequence $\hat{U}_{t}=\hat{U}_{t-1}^{*}$, reservoir state ${\mathbf{x}}_{t}$, output weight ${\mathbf{W}}^{\mathrm{out}}_{t}$, precision matrix ${\mathbf{P}}_{t}$, cost function $\ell$, $\phi$, inverse temperature $\lambda$, and noise covariances $\Sigma,\tilde{\Sigma}$.
$S_{k,{\tilde{k}}}\leftarrow 0,\ k\in\{0,\ldots,K-1\},\ {\tilde{k}}\in\{0,\ldots,{\tilde{K}}-1\}$
Sample $E=\{\left(\epsilon_{t}^{k},\ldots,\epsilon_{t+H-1}^{k}\right)\mid\epsilon_{\tau}^{k}\sim{\mathcal{N}}(0,\Sigma),\ k\in\{0,\ldots,{\tilde{K}}-1\},\ \tau\in\{t,\ldots,t+H\}\}$
Sample $W=\{\left({\mathbf{W}}_{t}^{{\tilde{k}}},\ldots,{\mathbf{W}}_{t+H}^{{\tilde{k}}}\right)\mid\mathrm{vec}({\mathbf{W}}^{{\tilde{k}}}_{\tau})\sim{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}^{\mathrm{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}),\ {\tilde{k}}\in\{0,\ldots,{\tilde{K}}-1\},\ \tau\in\{t,\ldots,t+H\}\}$
@@
    ${\mathbf{v}}^{k}_{\tau}\leftarrow\hat{\mathbf{u}}_{\tau}+\epsilon_{\tau}^{k}$
@@
       $S_{k,{\tilde{k}}}\mathrel{+}=\ell\left({\mathbf{W}}_{\tau}^{{\tilde{k}}\ \top}\hat{\mathbf{x}}^{k}_{\tau}\right)+\lambda(\hat{\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}})^{\top}\Sigma^{-1}({\mathbf{v}}^{k}_{\tau}-{\mathbf{u}}^{\text{ref}})$
@@
    $S_{k,{\tilde{k}}}\mathrel{+}=\phi\left({\mathbf{W}}_{H}^{{\tilde{k}}\ \top}\hat{\mathbf{x}}^{k}_{H}\right)$
@@
$\hat{U}^{*}_{t}\leftarrow\hat{U}_{t}+\tilde{\eta}^{-1}\sum_{k=0}^{K-1}\left[\sum_{{\tilde{k}}=0}^{{\tilde{K}}-1}\exp\left(-\lambda^{-1}S_{k,{\tilde{k}}}\right)\left(\epsilon_{t}^{k},\ldots,\epsilon_{t+H-1}^{k}\right)\right]$
@@
The resulting UMPPI algorithm is presented in Algorithm 3. In this algorithm, the expectation in 32 is approximated by Monte Carlo integration using $K$ samples of input noise and $\tilde{K}$} samples of output weight matrices, yielding an approximate optimal input sequence $\hat{U}^{*}_{t}$. To improve sampling efficiency, we use the input sequence from the previous control cycle, $\hat{U}_{t}=\hat{U}^{*}_{t-1}$, as the initial estimate of the input sequence $\hat{U}_{t}$. Since the input and output noise are independent random variables, Monte Carlo sampling can be efficiently executed in parallel by pre-generating batches of samples.

### v2

@@ IV-A Echo State Network
We describe the ESN model and its learning method. The ESN is a lightweight model capable of approximating complex nonlinear dynamics. We employ the ESN as a probabilistic model to quantify the uncertainty of the learned dynamics, and employ RLS [40, 41] for online learning of the model parameters.
IV-A1 Model and Learning Algorithm
As a surrogate model in 6 and 7, we employ the ESN with leaky terms, which incorporates temporal dynamics through leak rates of the target control system [20]:
$\begin{aligned} {\mathbf{x}}_{t+1}&=\hat{\mathbf{F}}({\mathbf{x}}_{t},{\mathbf{u}}_{t})\\ &\coloneqq(1-\alpha){\mathbf{x}}_{t}+\alpha\,f\left({\mathbf{W}}^{\text{res}}{\mathbf{x}}_{t}+{\mathbf{W}}^{\text{in }}{\mathbf{u}}_{t}\right),\end{aligned}$ (8)
@@ IV-A1 Model and Learning Algorithm
Here, $^{\mathbf{x}}_{t}\in{\mathbb{R}}^{\hat{N}}$ is the reservoir state vector, where $\hat{N}\in{\mathbb{N}}$, and $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{L}$ is the output vector. Additionally, $^{\mathbf{W}}^{\text{res}}\in{\mathbb{R}}^{\hat{N}\times\hat{N}}$ and ${\mathbf{W}}^{\text{in}}\in{\mathbb{R}}^{\hat{N}\times M}$ are the state and input weight matrices, respectively, and ${\mathbf{W}}^{\text{out}}_{t}\in{\mathbb{R}}^{\hat{N}\times L}$ is the output weight matrix. Furthermore, $f\colon{\mathbb{R}}^{\hat{N}}\to{\mathbb{R}}^{\hat{N}}$ is an element-wise nonlinear activation function, and $\alpha\in(0,1]$ is a parameter called the leak rate.
The ESN is a lightweight model that enables fast learning of dynamical systems with nonlinear characteristics. This computational efficiency arises because, unlike conventional NNs that train all weight matrices, only the output weight matrix ${\mathbf{W}}^{\text{out}}_{t}$ needs to be trained. The state and input weight matrices ${\mathbf{W}}^{\text{res}}$ and ${\mathbf{W}}^{\text{in}}$ remain fixed as randomly initialized matrices. The reservoir state dimension $\hat{N}$ is set greater than the input or output dimensions. The matrix ${\mathbf{W}}^{\text{res}}$ is a sparse random matrix with many components equal to zero, designed empirically so that the maximum absolute value of eigenvalues (spectral radius) is 1 or less. The matrix ${\mathbf{W}}^{\text{in}}$ is randomly generated from a uniform distribution. The reservoir state dimension $\hat{N}$, sparsity and spectral radius of ${\mathbf{W}}^{\text{res}}$, the shape of the activation function $f$, the leak rate $\alpha$, and the range of the input weight matrix ${\mathbf{W}}^{\text{in}}$ are hyperparameters that must be set prior to learning.
In online settings where data arrives sequentially, the ESN enables particularly rapid learning through RLS. This algorithm minimizes the following weighted least squares error:
@@
where $\gamma\in(0,1]$ is the discount factor. The following proposition summarizes the recursive update rule.
@@
Given : Discount factor $\gamma$, previous output weight ${\mathbf{W}}^{\text{out}}_{t-1}$, precision matrix ${\mathbf{P}}_{t-1}$, observed output ${\mathbf{y}}_{t}$, and current state ${\mathbf{x}}_{t}$.
return ${\mathbf{W}}^{\text{out}}_{t}$ and ${\mathbf{P}}_{t}$
Algorithm 2 RLS [40]
[…]
IV-A2 Probabilistic Interpretation
We treat the ESN as a probabilistic model to quantify the uncertainty of the model during learning. Given that ${\mathbf{W}}^{\text{out}}_{t}$ is the only learnable parameter in the ESN, we examine the probability distribution of ${\mathbf{W}}^{\text{out}}_{t}$ and provide a probabilistic interpretation of 8 and 9.
In RLS, ${\mathbf{W}}^{\text{out}}_{t}$ is updated sequentially with each newly observed data. In the early stages of learning, the estimation error is large due to the limited amount of observed data. As the amount of data becomes sufficiently large, the RLS estimate converges to the ordinary least squares estimator. Therefore, we define $\bar{\mathbf{W}}^{\text{out}}$ as the limit of ${\mathbf{W}}^{\text{out}}_{t}$ as $t\to\infty$ and define the learning error as follows:
$\mbox{$\epsilon$}_{t}\coloneqq{\mathbf{y}}_{t}-\bar{\mathbf{W}}^{\text{out}\top}{\mathbf{x}}_{t}.$ (13)
We assume that this error $\mbox{$\epsilon$}_{t}\in{\mathbb{R}}^{L}$ behaves as stochastic noise with mean 0 and covariance matrix $\tilde{\Sigma}=\operatorname*{diag}(\sigma_{1}^{2},\ldots,\sigma_{L}^{2})$ independently at each time. Although this assumption acts as an approximation for the complex error dynamics in practice, it allows us to derive the following probabilistic properties of ${\mathbf{W}}^{\text{out}}_{t}$.
@@ IV-A2 Probabilistic Interpretation
If $\gamma=1$,
${\mathbb{E}}[\mathrm{vec}({\mathbf{W}}^{\text{out}*}_{t}-\bar{\mathbf{W}}^{\text{out}})\mathrm{vec}({\mathbf{W}}^{\text{out}*}_{t}-\bar{\mathbf{W}}^{\text{out}})^{\top}]=\tilde{\Sigma}\otimes{\mathbf{P}}_{t}.$ (14)
This property allows us to quantify the model uncertainty $\mathcal{P}_{t}$ introduced in 5. By stochastically sampling the output weight matrix using ${\mathbf{P}}_{t}$, we can approximately evaluate the expected value in 5. Furthermore, since the ESN represents uncertainty solely through the output weight matrix, this sampling strategy provides computational advantages over methods that propagate uncertainty through the entire state vector (see Remark IV.8). Note that Proposition IV.2 is obtained by extending the analysis in Ref. [40] to the case of ESNs, where the estimated parameters take a matrix form; a detailed proof is provided in the Appendix.
@@ IV-B Uncertainty-aware Model Predictive Path Integral Control
We propose control methods combining the ESN with MPPI or its uncertainty-aware variant UMPPI. In MPPI [27], multiple noisy input sequences are sampled based on a reference input to simulate future trajectories of the control system, and an output cost is computed for each sequence. The optimal control input is approximately computed as a weighted average of many samples, where samples with lower costs receive higher weights. In UMPPI, in addition to sampling input sequences, output weight matrices of the ESN with noise added are sampled. This provides control input perturbations according to the identification accuracy, allowing the controller to account for model uncertainty and achieve robust performance. Note that in Algorithm 3 described below, when the sample size ${\tilde{K}}=1$ and the noise covariance $\tilde{\Sigma}=0$, UMPPI reduces to the standard MPPI algorithm. Therefore, we focus on the UMPPI in the following description.
@@
The following derivation of UMPPI extends the framework of Ref. [27] by introducing perturbations of the output-weight matrix. This extension preserves the mathematical consistency of the MPPI formulation and constitutes a theoretical contribution of the present work. Specifically, Proposition IV.4 parallels Eq. (13) in Ref. [27]; Proposition IV.5 is established by the same argument as in Section III-A of that reference; and Proposition IV.6 corresponds to Eq. (27) in the same reference.
We outline how UMPPI computes the optimal control solution through a sequence of sampling-based approximations. The derivation proceeds in the following three steps:
@@
    Stochastic reformulation. The stochastic optimal control problem in 5 is approximated by introducing noise into the control inputs and model parameters, as expressed in 21.
@@
    Free-energy relaxation. Introducing a free-energy functional converts this problem into the minimization of a Kullback–Leibler divergence between two probability distributions, as shown in 32.
@@
The details of each step are described in Sections IV-B1, IV-B2 and IV-B3.
@@ IV-B1 Stochastic reformulation
This subsection reformulates the original MPC problem in 5 as an optimization problem over perturbed control inputs and model parameters in 21. Like conventional MPPI, UMPPI considers perturbed inputs ${\mathbf{v}}_{t}\sim{\mathcal{N}}({\mathbf{u}}_{t},\Sigma)$ where Gaussian noise is applied to some initial inputs ${\mathbf{u}}_{t}$:
$\hat{\mathbf{x}}_{\tau+1}=\hat{\mathbf{F}}(\hat{\mathbf{x}}_{\tau},{\mathbf{v}}_{\tau}),\ \hat{\mathbf{x}}_{t}={\mathbf{x}}_{t},$ (15)
where $\hat{\mathbf{F}}$ is defined in 8. Furthermore, $\Sigma\in{\mathbb{R}}^{M\times M}$ represents the covariance matrix of the applied noise, and its value is determined based on the input weight matrix $R$ of the cost function (3):
@@
where $\lambda\in{\mathbb{R}}$ is a parameter called the inverse temperature, which is a design parameter. Additionally, UMPPI considers the output weight matrices ${\mathbf{W}}_{\tau}$ with noise applied to the reference weight matrix ${\mathbf{W}}^{\text{out}}_{\tau}$ at each time:
@@
where $\mathrm{vec}({\mathbf{W}}_{\tau})\sim{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}^{\text{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t})$, which is justified by Proposition IV.2. Applying noise to inputs enables approximate computation of optimal inputs through the relationship with free energy, as explained in the following subsections. Additionally, applying noise to the output weight matrix allows us to approximate the expectation ${\mathbb{E}}_{\mathcal{P}_{t}}[\cdot]$ in (5), which results in a control strategy robust to identification errors.
@@
$\begin{aligned} &p(W_{t}|{\mathbf{W}}_{t}^{\text{out}},{\mathbf{P}}_{t},\tilde{\Sigma})\\ &\quad=\prod_{\tau=t}^{t+H}{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}_{\tau})|\mathrm{vec}({\mathbf{W}}^{\text{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}).\end{aligned}$ (20)
@@ IV-B2 Free energy relaxation
This subsection introduces a variational relaxation of the stochastic optimization in 21 using the concept of free energy, leading to an alternative minimization problem in 32. We define the free energy $\mathcal{F}$} as follows:
@@
$q(W_{t})$ $\coloneqq\prod_{\tau=t}^{t+H}{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}_{\tau})|\mathrm{vec}({\mathbf{W}}^{\text{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}).$ (27)
Here, ${\mathbf{u}}^{\text{ref}}$ is the reference input defined in 3. The following two propositions show the relationship between LABEL:eq:free_energy and 21.
@@
Suppose that $V_{t}$ and $W_{t}$ in 19 and LABEL:eq:prob_W follow the distributions given by
@@
Then, equality holds for 28:
@@ IV-B3 Sampling-based optimization
Suppose that the admissible control input set is given by ${\mathcal{U}}={\mathbb{R}}^{M\times H}$. Then, given any input sequence $\hat{U}_{t}$, the control input satisfying 32 is computed as
@@
Given : Sample size of input/output perturbation $K,{\tilde{K}}$, prediction horizon $H$, initial estimate of input sequence $\hat{U}_{t}=\hat{U}_{t-1}^{*}$, reservoir state ${\mathbf{x}}_{t}$, output weight ${\mathbf{W}}^{\text{out}}_{t}$, precision matrix ${\mathbf{P}}_{t}$, cost function $\ell$, $\phi$, inverse temperature $\lambda$, and noise covariances $\Sigma,\tilde{\Sigma}$.
$J_{k,{\tilde{k}}}\leftarrow 0,\ k\in\{0,\ldots,K-1\},\ {\tilde{k}}\in\{0,\ldots,{\tilde{K}}-1\}$
Sample $E=\{\left({\mathbf{e}}_{t}^{k},\ldots,{\mathbf{e}}_{t+H-1}^{k}\right)\mid{\mathbf{e}}_{\tau}^{k}\sim{\mathcal{N}}(0,\Sigma),\ k\in\{0,\ldots,{\tilde{K}}-1\},\ \tau\in\{t,\ldots,t+H\}\}$
Sample $W=\{\left({\mathbf{W}}_{t}^{{\tilde{k}}},\ldots,{\mathbf{W}}_{t+H}^{{\tilde{k}}}\right)\mid\mathrm{vec}({\mathbf{W}}^{{\tilde{k}}}_{\tau})\sim{\mathcal{N}}(\mathrm{vec}({\mathbf{W}}^{\text{out}}_{t}),\tilde{\Sigma}\otimes{\mathbf{P}}_{t}),\ {\tilde{k}}\in\{0,\ldots,{\tilde{K}}-1\},\ \tau\in\{t,\ldots,t+H\}\}$
@@
    ${\mathbf{v}}^{k}_{\tau}\leftarrow\hat{\mathbf{u}}_{\tau}+{\mathbf{e}}_{\tau}^{k}$
@@
       $J_{k,{\tilde{k}}}\mathrel{+}=\ell\left({\mathbf{W}}_{\tau}^{{\tilde{k}}\ \top}\hat{\mathbf{x}}^{k}_{\tau}\right)+\lambda(\hat{\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}})^{\top}\Sigma^{-1}({\mathbf{v}}^{k}_{\tau}-{\mathbf{u}}^{\text{ref}})$
@@
    $J_{k,{\tilde{k}}}\mathrel{+}=\phi\left({\mathbf{W}}_{H}^{{\tilde{k}}\ \top}\hat{\mathbf{x}}^{k}_{H}\right)$
@@
$\hat{U}^{*}_{t}\leftarrow\hat{U}_{t}+\tilde{\eta}^{-1}\sum_{k=0}^{K-1}\left[\sum_{{\tilde{k}}=0}^{{\tilde{K}}-1}\exp\left(-\lambda^{-1}J_{k,{\tilde{k}}}\right)\left({\mathbf{e}}_{t}^{k},\ldots,{\mathbf{e}}_{t+H-1}^{k}\right)\right]$
@@
The resulting UMPPI algorithm is presented in Algorithm 3. In this algorithm, the expectation in 33 is approximated by Monte Carlo integration using $K$ samples of input noise and $\tilde{K}$} samples of output weight matrices, yielding an approximate optimal input sequence $\hat{U}^{*}_{t}$. To improve sampling efficiency, we use the input sequence from the previous control cycle, $\hat{U}_{t}=\hat{U}^{*}_{t-1}$, as the initial estimate of the input sequence $\hat{U}_{t}$. Since the input and output noise are independent random variables, Monte Carlo sampling can be efficiently executed in parallel by pre-generating batches of samples.